Three interior angles of a pentagon are each 3y and the remaining each twice the first three. Find the value of y. Please help me solve this
Answer by ikleyn(52792)

.
Three interior angles of a pentagon are each 3y and the remaining each twice the first three.
Find the value of y. Please help me solve this
~~~~~~~~~~~~~~~~

The sum of interior angles of any n-gon is  (n-2)*180°.


For a pentagon, it is  (5-2)*180° = 3*180° = 540°.


The three angles are 3y each;  the remaining two angles are 2*(3y) = 6y each.


An equation to find y is

    3y + 3y + 3y + 6y + 6y = 540°,

or

    21y = 540  degrees,


which implies

    y = 540/21 = 180/7.


ANSWER.  The value of y is  180/7 degrees.

Solved.


////////////////


Had the author be willing to create an educationally valuable problem,
he (or she) would take care to pick up more rounded numbers.